OpenStudy (anonymous):

Sin{arcsin(1/2)+arccos(5/12)=?

OpenStudy (anonymous):

first convert arccos(5/12) in terms of arcsin(?)

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

let arccos(5/12) = x then cos x =5/12 and now find sin x

OpenStudy (anonymous):

\[\sqrt{119}/12\]

OpenStudy (anonymous):

my answer to this prob is \[(5+\sqrt{357})/24\]

OpenStudy (anonymous):

is it correct?